A skydriver of mass 80kg falls vertically with a constant speed of 50 m/s. The upward force acting on the skydriver is approximately
A.0N B.80N C.800N D 4000N
Answer #1

Solution:

Since the sky diver falls vertically with a constant speed, it's downward acceleration is zero. Since net force on the sky diver, F=M*a, where M is the mass of the sky diver and a is the net acceleration,

Net force F=80*0 N =0 N M=80 kg

Since the net force on the sky diver is zero, the upward force must cancel off the downward force due to gravity, F'
F'=M*g = 80*10 N =800 N. Where g is the acceleration due to gravity = g = 10 m/s^2

Hence the upward force on the sky diver is 800 N or option C.

Note: The actual value of g is 9.8 m/s^2 but the approximate value used here is 10m/s^2.